absolute value
|5x-2|+4= 7
|5x-2| =7-4
tx-2=3
5x=3+2
x=1/5 times 5
x=1
Now the second part is what I am not getting
|5x-2|+4=-7
5x-2=-7-4
5x=-11+2
5x=-9
This is not right...what am I doing wrong???

Answer by ptth176(50)

You wrong
|5x-2|+4= 7
=>|5x-2|=3
=>5x-2=3 or 5x-2=-3
=>x=1 or x=-1/5